OpenStudy (anonymous):

A soccer goalie kicks a ball. The ball leaves the goalie's foot with an initial upward velocity of 10 meters per second. If the ball is 1.5 meters off the ground when it leaves the goalie's foot, about how many seconds will it take for the ball to hit the ground?

OpenStudy (ledah):

whoa o_o

OpenStudy (ledah):

do you know the formula?

OpenStudy (anonymous):

I know right :/

OpenStudy (anonymous):

no I dont

OpenStudy (ledah):

we can use this one V = (P2-P1)/T or this one i believe is another one v=d/t

OpenStudy (anonymous):

Its multiple choice the possible answers are...

OpenStudy (anonymous):

A=3/2b^2 A=9/2B^2 A=1/6b^2 A=1/18b^2 A = 3 b2 2

OpenStudy (anonymous):

disregard anything except the 1st 4

OpenStudy (ledah):

This is a difficult one.

OpenStudy (ledah):

you could try the physics section, see if that subject has better help. because this is pretty difficult. I dont think I can help you with this. I apologize

OpenStudy (anonymous):

Okay thanks for trying\

OpenStudy (ledah):

no problem. I'm sorry I cant do much more for you.